OpenStudy (anonymous):

5. True or false. P(Ec ∪ Fc) = 1 - P(E ∩ F)

OpenStudy (amistre64):

Ec mean compliment of E ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then (E' u F')' = (E n F)

OpenStudy (amistre64):

A' = 1 - A (E' u F')' = 1 - (E' u F') = 1 - (E n F)' hmm

OpenStudy (amistre64):

oy my head hurts!!

OpenStudy (amistre64):

P(E' u F') P(E n F)' = 1 - P (E n F)

OpenStudy (amistre64):

sooks good to me; we can run a venn on it to be sure :)

OpenStudy (amistre64):